# LENVIZ: A High-Resolution Low-Exposure Night Vision Benchmark Dataset

### v1

@@ III LENVIZ Capture
Our dataset is composed of $24,082$ low-light scenes captured over $3$ different Complimentary Metal-Oxide-Semiconductor (CMOS) image sensors. We aimed to create a diverse dataset suitable for pixel-level image processing tasks such as image enhancement and high dynamic range (HDR) imaging in challenging low-light conditions. With this in mind, we provide up to 9 multi-exposure frames as well as a long-exposure shot for each scene ($234,688$ total frames) to aid in the development of single, and multi-exposure image enhancement approaches. In this section we detail the capture pipeline for our data as well as sensor-specific details and configurations.
III-A Image Sensor Information
To broaden our dataset’s representativeness and improve the generalizability of algorithms trained using our data, we provide scenes captured over $3$ different non-DSLR sensors. Information on the specific details (aperture, sensor-size, FOV, etc.) of each is provided in tabl III. While our S5K4H7YX03-FGX9 Sensor is better suited for close range captures, S5KJN1SQ03 and S5KJNS camera features (showcased in Table III) allow for richer capture of details at medium and long range.
TABLE III: Camera Module Specifications
Sensor Name
Resolution
Aperture
Pixel-size
Sensor-Size Focus FOV (Diag)
S5K4H7YX03-FGX9 8 f/2.0 1.12 1/3 Fixed (27$\sim$39.3cm) 78°
[…]
It is important to note that, the performance and characteristics of sensors can vary based on their placement on a device and intended function, making certain sensors more suitable for specific applications. We used mobile devices to capture our dataset, for these there are typically two main cameras: one on the front and one on the rear. Our first sensor, S5K4H7YX03-FGX9, was the front-facing camera on the capture device. These cameras are essential for selfies and video calls but are constrained in physical size due to their placement above the phone display. Although their resolution is generally lower than that of rear cameras, they are designed to capture scenes at close range (usually an arm’s length) and often benefit from additional light from the screen. As such, our images from this sensor tend to contain a higher concentration of faces Fig. 7.
The other two sensors, S5KJN1SQ03 and S5KJNS, were used as rear cameras on different devices. Rear cameras are designed to serve as the primary lens for capturing the environment. Unlike front cameras, they offer higher resolution and specifications for more detailed scenes, which typically involve greater depth and lower lighting conditions. While face detection remains important for rear cameras, they are also used for capturing a wider range of subjects, such as landscapes and objects.
Figure 2: [Top] Lux vs exposure time (log scale). The graph is plotted for standard values of ISO 100, aperture 1.8 and no additional boost. [Bottom] Lux vs exposure time (log scale) for long-exposure ground truth image.
III-B Ambient Light Lux Calculation
For our data acquisition, we implemented stringent controls over camera parameters such as ISO sensitivity, exposure time, and aperture. Concurrently, we performed an estimation of the scene lux levels using our capturing devices, this measurement indicates the intensity of light in a scene and is essential for characterizing low-light environments and is included for each scene. This deliberate approach ensured a highly controlled environment for image capture, thereby enhancing the dataset’s suitability for benchmarking low-light image enhancement algorithms.
We estimate the lux readings of the captured images using key camera parameters like exposure time, ISO sensitivity, and lens aperture. Our lux calculation is as follows:
$\text{lux}=\beta_{\text{lux}}\times f^{2}\times\left(\frac{10^{11}}{t_{\text{exp}}\times ISO}\right)$ (1)
Here, the exposure time ($t_{\text{exp}}$) determines the duration (in $ns$) that the camera sensor is exposed to light. ISO sensitivity ($ISO$) controls the sensor’s sensitivity to light, with higher values increasing sensitivity but also introducing noise. The lens aperture ($f$) regulates the amount of light reaching the sensor. The f-value is fixed for all sensors in our dataset. Finally, camera-specific tuning constants ($\beta_{\text{lux}}$) account for sensor characteristics and ISP tuning.
As we can see in Equation 1, our lux estimation is inversely proportional to both the exposure time and the ISO value. Longer exposure times and higher ISO values reduce the calculated lux. The aperture size $f$, directly influences the light-gathering capacity of the lens. Additionally, when a post-processing sensitivity boost is applied (e.g., through ISO adjustment as part of the ISP tuning of a sensor), the ISO value is modified by the post-boost percentage. The estimated lux is plotted against exposure time for standard values of ISO, aperture and boost in Fig. 2 [Top].
III-C Exposure Time Calculation for Long-Exposure Shots
The exposure time for our long-exposure images is computed based on the ambient light intensity, measured in lux, using a formula derived from the relationship between lux and exposure time. Equation 2 governs the exposure time calculation for the long-exposure ground truth image given the estimated lux of the scene as follows:
[…]
In scenarios where the lux value is less than 8, the exposure for long-exposure image is determined as $\frac{48}{\text{lux}}$. For lux values between 8 and 10, the exposure is slightly modified using a power function, $\frac{48}{\text{lux}^{0.99}}$, to ensure a smoother transition between lighting conditions. When the lux value falls between 10 and 15, the exposure increases to $\frac{52}{\text{lux}}$, and for values above 15, the exposure is further extended to $\frac{60}{\text{lux}}$.The estimated exposure time for long-exposure ground truth image is plotted against the estimated lux in Fig. 2 [Bottom].
III-D Post Capture Pipeline
@@
Existing benchmark datasets for low-light image enhancement are primarily composed of either RAW[26, 7, 9] or JPEG[18, 1, 16] images. A contrast between the traditional camera sensor-to-gallery pipeline and ours is demonstrated in Fig. 3. While RAW images offer a more pristine representation of the captured scene, their use is often restricted to commercial devices that provide access to raw data. JPEG images, on the other hand, are typically gallery images that have undergone extensive camera ISP tuning and default post-processing (such as low-light face beautification, noise reduction, over-exposure suppression, etc). This processing can significantly alter the original sensor data, leading to a loss of fine details. As a result, models trained on these datasets may struggle to learn how to deal with sensor/ISP-tuning specific noise and fail to generalize to camera images where the camera’s ISP and post-processing algorithms may differ.
@@
LENVIZ addresses the limitations of existing datasets by providing a collection of JPEG images that are exclusively ISP-tuned, devoid of any default camera post-processing. This approach ensures that the data closely resembles the raw output of the sensor+ISP pipeline, capturing the inherent noise, over-exposure, and other artifacts that are characteristic of low-light conditions. By training enhancement models on these ISP-tuned-only input frames, we effectively leverage the expertise of state-of-the-art ISP processing while also enhancing learning efficiency.
IV LENVIZ Properties
@@
In the following sections we will detail the specifics on the LENVIZ data distribution across different luminance levels, capture scenarios and content classes. Table II provides a detailed breakdown of the number of frames, scenes, and human-generated ground-truth in our dataset.
IV-A Luminance conditions
We adopted lux ranges as one of the primary criteria for selecting low-light scenes. Lux distribution plays a crucial role in building datasets for low-light image capture because it provides a quantitative measure of the scene’s overall illumination. We ensured that the dataset encompasses a variety of lighting conditions, from dimly lit rooms to outdoor scenes illuminated by street-lights.
@@ IV-A Luminance conditions
Our dataset was captured over a wide range of lux-levels, ranging from extreme low-light (lux $\approx$ 0), to brighter scenes captured at dusk or illuminated by varying light sources. Fig. 4 illustrates the lux distribution of the data captured across all 3 sensors. We paid special interest to the extreme-low light ranges to address the main shortcomings that most of the benchmark datasets are lacking. Images captured at such low levels of illumination provide a more challenging problem as they are more severely impacted by noise, lack of details, and color information and greatly benefit from having high quality annotations to serve as a guide for deep learning models on how to address each of these issues.
Figure 4: LENVIZ Lux Histogram
IV-B Capture scenarios
@@
One of the key aspects of LENVIZ is the richness of data we capture, demonstrated by our wide variety of scene types (see Fig. 1). We have captured indoor locations ranging from regular living/dwelling areas, specialized photography studios, gaming and entertainment rooms, conference auditoriums, and classroom setups among the most notable ones; and outdoor scenes containing both urban and rural locations showcasing a wide range of backgrounds like general city buildings and night lights, greenery, fountains and other water sources, etc. We made sure to provide a variety of illuminations sources with different intensities, colors, and complexities. These manifest in the forms of LED lights and signage both indoor and outdoors, string lights, directional lights, natural light, and ambient illumination.
@@
Indoor scenes: To comprehensively evaluate low-light image enhancement algorithms, our dataset includes variety indoor scenes characterized by uneven lighting and mixed light sources. This encompasses scenes with contrasting levels of illumination, such as brightly lit areas juxtaposed with dark corners or hallways. Additionally, we have captured scenes featuring a combination of incandescent and fluorescent lighting, whose differing color temperatures can pose significant challenges for image processing. These provide a robust testing ground for assessing the performance and generalizability of enhancement algorithms under various lighting conditions.
@@
Outdoor scenes: As a complement to our indoor scenes, our dataset also includes a variety of outdoor scenes designed to challenge the capabilities of low-light image enhancement algorithms. These scenes cover scenarios with high dynamic ranges, where there’s a stark contrast between brightly lit areas and darker elements. Additionally, we have captured scenes with strong directional lighting, such as sunsets or streetlamps, which introduce pronounced shadows. These outdoor scenarios provide a complex set of challenges for training and evaluating the performance of enhancement algorithms in real-world conditions.
IV-C Content classes
Figure 5: 14 Skin-tone variations and hyper-realistic mannequins used in the LENVIZ dataset replicating the most real-world human subject specific scenarios.
@@ IV-C Content classes
LENVIZ includes a wide variety of subjects and incorporates generous variations in scene content. We paid special attention to the presence of human and human-like subjects in our data, and to better encompass real-world conditions we incorporated a wide diversity of subject skin tones (14 mannequins spanning a diverse range of skin tones and facial features to ensure inclusivity), overall scene appearance (e.g., indoor vs. outdoor environments), and subject position (relative to the camera and light source). Since our long-exposure capture requires a higher restriction in motion, we limited the presence of humans to our test dataset, our train dataset contains 9 realistic and 4 hyper-realistic (Fig. 5) mannequins instead of humans for added control in the scenes. These mannequins were meticulously selected to exhibit natural variations in skin-tone, facial textures, pose, expression (simulated through subtle head tilts and eye positioning), and accessories (such as glasses). This approach introduces a high degree of realism while maintaining control over lighting conditions and scene composition, ultimately leading to a robust and generalizable dataset for low-light image processing tasks. Fig 7 presents a histogram of the distribution of human-in-scene data across three sensor modules. The results demonstrate a significantly larger distribution of face images captured by the front camera sensor (S5K4H7YX03-FGX9), reflecting the common selfie-style usage pattern. Overall, $70$% of our scenes contain one or more faces, whose variation enhances the model’s ability to learn from diverse camera usage scenarios.
@@
To further characterize the dataset’s scene content, we conducted a comprehensive object and face detection analysis using the AWS Label Detection [^4] https://docs.aws.amazon.com/rekognition/ tool. This analysis served two primary purposes: 1) assessing the overall capability of commercial approaches to detect objects in extreme low-lux conditions, and 2) to identify unique features and content richness within each scene. The results of this analysis were incorporated as metadata for each of the scenes of our dataset, including detected bounding boxes, object classes (230 object labels belonging to 27 distinct categories), and in the case of face detection, face descriptors such as gender, presence of facial hair, or glasses.
Fig 6 provides an overlook at the distribution of the different object categories identified by the AWS object detection algorithm. In general, the detected objects within our images can be classified into 27 broad categories, ranging from ”Plants” and ”Animals” to ”Buildings” and ”Vehicles”. Each of these categories are further divided into 230 object-specific labels like ”Chair” or ”Couch” for the general ”Furniture” Category. Appendix A Table VI provides a breakdown of each of the 230 object labels and their respective categories. We include the list of object labels identified for each given scene as well as their bounding boxes in our dataset release.
IV-D Feature Embedding Analysis
Figure 7: Number of Faces Per Scene Histogram
To understand the underlying structure and relationships within our LENVIZ dataset, we analyzed the feature distribution of our scenes in a latent space representation and compared it against other benchmakring datasets in the field, namely, LOL [1] (Single-Exposure), and SICE [12] (Multi-Exposure). We extracted a deep feature representation for each of the scenes of all three datasets using the output of the last convolutional block in the VGG16 model given its well known suitability for feature extraction and its ability to recognize a vast range of visual patterns. To reduce the dimensionality of these features and visualize their distribution, we applied t-SNE to map these high-dimensional data points to a lower-dimensional space while preserving their local structure. As seen in Fig 8, the t-SNE visualization reveals that both LOL and SICE have some degree of overlap. The distribution of features from the LOL datasets seems to be relatively widespread in comparison to SICE. Despite this, our dataset is able to encompass a much wider feature distribution than these two benchmark datasets, as such allowing for models to learn from a wider set of features when using LENVIZ.
IV-E Test Dataset
To allow for an extensive evaluation of image quality in low-light conditions, our test dataset was meticulously designed to encompass a wide range of parameters, including sharpness, texture, contrast, brightness, naturalness, accurate skin-tone, over-exposure recovery, and more. Drawing inspiration from the training dataset, we categorized scenes into two broad classes: indoor and outdoor. To account for diverse usage scenarios, we further differentiated between front and rear camera sensors, considering both selfie and non-selfie perspectives. Additionally, we incorporated human and non-human subjects in both indoor and outdoor settings to simulate real-world conditions. This resulted in a test dataset composed of 1,468 frames from 203 unique scenes. We provide two partitions: a “Reference” partition including a paired human-edited ground truth and long-exposure shot, and a “No-reference” partition which includes handheld captures and human subjects for a more challenging evaluation. The basic statistics of test dataset are provided in Table IV.
To introduce illumination variations, we systematically adjusted lux levels from 0 to 20 for rear imagers (S5KJN1SQ03 & S5KJNS) and from 0 to 30 for front imagers (S5K4H7YX03-FGX9). Furthermore, we included both flash-on and flash-off conditions to replicate authentic low-light scenarios encountered in everyday photography.
TABLE IV: Summary of the LENVIZ Test Dataset.
@@ IV-E Test Dataset
V LENVIZ Application
TABLE V: Quantitative Evaluation results for SOTA single exposure (SE) methods and Multi-exposure fusion (ME) methods on LENVIZ dataset. Here, higher the PSNR and SSIM value and lower the LPIPS score indicated that better the output quality
Type SOTA Test Dataset $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS
LENVIZ
trained
LOL/SICE
trained
LENVIZ
trained
LOL/SICE
trained
LENVIZ
trained
LOL/SICE
trained
SE ZeroDCE++ [62] LENVIZ 16.35 16.38 0.359 0.437 0.663 0.611
LOL 14.75 14.75 0.518 0.518 0.328 0.328
LLFormer [18] LENVIZ 21.13 17.05 0.665 0.551 0.358 0.498
LOL 19.33 19.79 0.757 0.772 0.240 0.277
ExpoMamba [63] LENVIZ 19.8 17.04 0.585 0.534 0.524 0.599
LOL 23.65 18.55 0.816 0.759 0.169 0.291
MEFNet [67] LENVIZ 20.71 20.77 0.609 0.606 0.457 0.458
SICE 21.13 20.94 0.612 0.612 0.358 0.361
HoLoCo [66] LENVIZ 21.31 20.77 0.613 0.606 0.689 0.689
SICE 13.78 13.90 0.614 0.615 0.526 0.529
MobileMEF [68] LENVIZ 20.93 19.47 0.626 0.613 0.492 0.561
SICE 13.65 14.36 0.637 0.632 0.484 0.452
To fully showcase the usability of our dataset for different approaches for low-light enhancement, we provide an analysis on $3$ Single Exposure and $3$ Multi Exposure low-light enhancement methods. Single exposure enhancement methods often use a low-exposure frame and attempt to improve its brightness, color, and detail recovery while reducing noise. Multi-exposure methods on the other hand incorporate the information of multiple frames captured at different exposure values, this often involves the use of a low exposure frame to recover information of over-exposed regions, and a medium or high exposure frame to get information related to color and details in under-exposed regions of the scene. All the $6$ approaches were trained using the LENVIZ dataset, as well as state-of-the-art single and multi-exposure benchmark datasets, LOL [1], and SICE [12] respectively.
Single Exposure Approaches: Zero-DCE++ [62], leverages a zero-reference deep curve estimation technique to enhance images without the need for paired ground-truth data, focusing on real-time illumination adjustment to improve brightness and contrast. LLFormer [60], incorporates axis-based multi-head self-attention and cross-layer attention fusion. Finally, ExpoMamba [63] introduces a novel architecture that integrates frequency state space components in a Mamba (a state space model family) to tackle real-time processing challenges.
Multi-Exposure Approaches: MEFNet [67] employs a multi-exposure fusion network that predicts the fusion weight maps at a low resolution for fast processing, HoLoCo [66] introduces contrastive learning with a holistic and local contrastive constraint to multi-exposure image fusion to recover details and allow for uniform illumination in over and under exposed regions. MobileMEF [68] is a lightweight multi-exposure fusion network for real-time processing on mobile platforms.
Figure 10: Representing model outputs for the state-of-the-art low light model performances trained on LENVIZ dataset.
V-A Quantitative Analysis
The performance of the evaluated models is summarized in Fig. 10 and Table V. The quantitative analysis was performed on the LENVIZ Reference Test data ($60$ scenes), and $15$ scenes from both LOL, and SICE ($90$ images total). In the further section, LENVIZ-trained indicates that the model was trained using LENVIZ dataset and LOL/SICE trained refers to models that were trained from scratch using either of the benchmark datasets (LOL for single exposure methods, and SICE for multi-exposure methods).
We have observed that the models trained on LENVIZ data for the most part deliver higher performance across all three metrics. In cases where the models trained on LENVIZ data obtained lower overall scores than their LOL/SICE trained counterparts, the differences were non-significant as is the case for ZeroDCE++ [62], this could be attributed to its unsupervised training paradigm as it prevents it from leveraging high-quality ground truth data. Consequently, we can observe that ZeroDCE++ demonstrated slightly weaker performance compared to other supervised single-exposure methods.
(a) Single exposure models
(b) Multi exposure models
@@ V-A Quantitative Analysis
Figure 11: Comparison of model outputs trained on the LENVIZ and benchmark datasets. The histogram illustrates user preferences for model outputs from LENVIZ compared to others, while the spider chart highlights the advantages of various image quality aspects based on user votes.
In certain instances, PSNR values for LENVIZ-trained models were slightly lower (eg. MEFNet [67], HoLoCo [66], ZeroDCE++, etc) than those of benchmark-trained models. However, these differences were minimal, likely due to dataset disparities. Many benchmark-trained models rely on a narrow range of data, often neglecting real-world camera and ISP-tuned noise patterns. When confronted with highly variable input data, these models may require extended training times to converge, especially when dealing with large datasets like LENVIZ. Given our short training time (100 epochs), these models may have faced challenges in adapting to the diverse noise patterns present in real-world images while learning to improve the overall image quality. While PSNR has illustrated high fluctuations in the scores, LENVIZ-trained models consistently achieved higher LPIPS scores compared to the benchmark-trained counterparts. LPIPS [65], a perceptual similarity metric, aligns more closely with human perception, making it a valuable tool for evaluating image enhancement performance. Therefore, considering LPIPS scores and overall performance, LENVIZ-trained models demonstrate superior intelligence and improved stability due to their exposure to a wider range of real-world camera data variations, outperforming traditional benchmark-trained models.
V-B Qualitative Analysis
The composition and quality of a dataset are pivotal in determining the performance of computer vision models, as they directly impact the model’s ability to generalize and function effectively in real-world scenarios. Thus, curating a balanced and representative dataset is critical for producing robust and accurate model outcomes [64].
We conducted a thorough qualitative evaluation of state-of-the-art (SOTA) low-light enhancement models to assess the influence of the LENVIZ dataset on image quality. A pairwise human evaluation (A/B testing) was employed, where participants compared outputs from the two model training strategies LENVIZ-trained and LOL and SICE-trained. Participants were asked to express their preference between the pairwise outputs and provide the primary reason for their choice.
For a comprehensive evaluation, we selected 10 diverse scenes containing various objects, landscapes, and both human subjects and mannequins. The illuminance levels in these scenes ranged from 0.02 to 20 lux. A total of 238 participants were randomly chosen from a pool of 89,883 Prolific candidates [^6] https://www.prolific.com/, ensuring diverse perspectives and unbiased feedback on model performance.
To evaluate the performance of existing models on the LENVIZ dataset, we tested $6$ SOTA models for low-light image enhancement, as illustrated in Fig. 11. The models included three single-exposure models (LLFormer, ExpoMamba, ZeroDCE++) and three multi-exposure models (MEFNet, HoLoCo, MobileMEF). Pairwise comparisons were conducted for each model, contrasting outputs from models trained on the two dataset strategies. A user survey was then used to rank the images and identify the primary factors influencing their preferences.

### v2

@@ 3.3 Long-Exposure Time Calculation
3.4 Post Capture Pipeline
@@
Existing benchmark datasets for low-light image enhancement are primarily composed of either RAW[MIT5K, SID, VE-LOL] or JPEG[UHDLOL2023, LOL, DarkZurisch] images. A contrast between the traditional camera module-to-gallery pipeline and ours is demonstrated in Fig. 1. While RAW images offer a more pristine representation of the captured scene, their use is often restricted to commercial devices that provide access to raw data. JPEG images, on the other hand, are typically gallery images that have undergone extensive camera ISP tuning and default post-processing (such as face beautification, noise reduction, over-exposure suppression, etc). This processing can significantly alter the original camera module data, leading to a loss of fine details. As a result, models trained on these datasets may struggle to learn how to deal with camera module/ISP-tuning specific noise and fail to generalize to changes in camera’s ISP and post-processing.
@@
LENVIZ addresses the limitations of existing datasets by providing a collection of JPEG images that are exclusively ISP-tuned, devoid of any default camera post-processing. This approach ensures that the data closely resembles the raw output of the camera module+ISP pipeline, capturing the inherent noise, over-exposure, and other artifacts that are characteristic of low-light conditions. By training enhancement models on these ISP-tuned-only input frames, we effectively leverage the expertise of state-of-the-art ISP processing while also enhancing learning efficiency.
3.5 Human Curated Ground truth Generation
Our human ground truth comprises a set of $13,067$ scenes edited by a team of 7 expert photographers and editors. The process followed a standardized workflow and style guide. The editors used the long-exposure shot as the base for the ground truth, as it offered a high-brightness view of the scene with considerably less noise than the multi-exposure frames. They then corrected and restored details lost to over-exposure using the multi-exposure frames, and were instructed to match the mid-exposure frames’ color tones, focusing on technical quality: brightness, contrast and noise reduction. Each edited image was then submitted for quality control and reviewed by three expert camera scientists, who evaluated it against a checklist of image quality aspects including noise levels, sharpness, and the absence of common artifacts like ghosting or color fringing, in addition to assessing color fidelity. Only images that received unanimous approval across all metrics were designated as the final “Human GT” for the dataset. This structured and peer-reviewed approach provides a reproducible target for algorithms to learn from and ensures a high level of consistency and quality, providing a robust benchmark for training and evaluating low-light enhancement models. While the final product represents a specific, expertly-curated style, the reproducibility and rigor of our process make it an effective and reliable target for algorithm development.
4 LENVIZ Properties
@@
In the following sections we will detail the specifics on the LENVIZ data distribution across different luminance levels, capture scenarios and content classes. Table 2 provides a breakdown of the number of frames, scenes, and human-generated ground-truth in our dataset. Additional analysis such as feature distribution is provided in Supp 7.2.3
4.1 Luminance conditions
@@
Illuminance level is one of the primary criteria for selecting low-light scenes as it provides a quantitative measure of the scene’s overall illumination (crucial in low-light images). We ensured that LENVIZ encompasses a variety of lighting conditions, from dimly lit rooms to outdoor scenes illuminated by street-lights, and paid special interest to the extreme-low light ranges (details on the illuminance distribution of our data are provided in Supp 7.2.2). Images captured at such low levels of illumination provide a more challenging problem as they are more severely impacted by noise, lack of details, and color information and greatly benefit from having high quality annotations to serve as a guide for deep learning models on how to address these issues.
4.2 Capture scenarios
@@
One of the key aspects of LENVIZ is the richness of data we capture, demonstrated by our scene coverage (see Fig. LABEL:fig:teaser). We have captured indoor locations ranging from regular living/dwelling areas, specialized photography studios, gaming and entertainment rooms, conference auditoriums, and classroom setups among the most notable ones; and outdoor scenes containing both urban and rural locations showcasing a wide range of backgrounds like general city buildings and night lights, greenery, fountains and other water sources, etc. We made sure to provide a variety of illuminations sources with different intensities, colors, and complexities. These manifest in the forms of LED lights and signage both indoor and outdoors, string lights, directional lights, natural light, and ambient illumination.
@@
Indoor scenes: To comprehensively evaluate low-light image enhancement algorithms, our dataset includes indoor scenes characterized by uneven lighting and mixed light sources. This encompasses scenes with contrasting levels of illumination, such as brightly lit areas juxtaposed with dark corners or hallways. Additionally, we have captured scenes featuring a combination of incandescent and fluorescent lighting, whose differing color temperatures can pose significant challenges for image processing. These provide a robust testing ground for assessing the performance and generalizability of enhancement algorithms.
@@
Outdoor scenes: As a complement to our indoor scenes, our dataset also includes a variety of outdoor scenes designed to challenge the capabilities of low-light image enhancement algorithms. These scenes cover scenarios with high dynamic ranges, where there’s a stark contrast between brightly lit areas and darker elements. We have also captured scenes with strong directional lighting, such as sunsets or streetlamps, which introduce pronounced shadows. These outdoor scenarios provide a complex set of challenges for training and evaluating the performance of enhancement algorithms in real-world conditions.
4.3 Content classes
Figure 2: Skin-tone variations and hyper-realistic mannequins.
@@ 4.3 Content classes
LENVIZ covers a broad spectrum of subjects and incorporates generous variations in scene content. We paid special attention to the presence of human and human-like subjects in our data, and to better encompass real-world conditions we incorporated a rich array of subject skin tones (14 mannequins spanning a diverse range of skin tones and facial features to ensure inclusivity), overall scene appearance (e.g., indoor vs. outdoor environments), and subject position (relative to the camera and light source). Since our long-exposure capture requires a higher restriction in motion, we limited the presence of humans to our test dataset, our train dataset contains 9 realistic and 4 hyper-realistic (Fig. 2) mannequins instead of humans for added control in the scenes. These were meticulously selected to exhibit natural variations in skin-tone, facial textures, pose, expression (simulated through head tilts and eye positioning), and accessories (such as glasses). This approach introduces a high degree of realism while maintaining control over lighting conditions and scene composition, ultimately leading to a robust and generalizable dataset for low-light image processing tasks. Overall, 7070% of our scenes contain one or more faces, whose variation enhances the model’s ability to learn from diverse camera usage scenarios.
@@
To further characterize the dataset’s scene content, we conducted a comprehensive object and face detection analysis using the AWS Label Detection [^3] https://docs.aws.amazon.com/rekognition/ tool. This analysis served two primary purposes: 1) assessing the overall capability of commercial approaches to detect objects in extreme low-illuminance conditions, and 2) to identify unique features and content richness within each scene. The results of this analysis were incorporated as metadata for each of the scenes of our dataset, including detected bounding boxes, object classes (230 object labels belonging to 27 distinct categories), and in the case of face detection, face descriptors such as gender, presence of facial hair, or glasses. For a detailed breakdown on this please consult Supp. 7.2.
4.4 Test Dataset
Table 3: Summary of the LENVIZ Test Dataset.
@@ 4.4 Test Dataset
To allow for an extensive evaluation of image quality in low-light conditions, our test dataset was designed to encompass a range of parameters, including sharpness, texture, contrast, brightness, naturalness, accurate skin-tone, over-exposure recovery, and more. This resulted in a test dataset composed of 1,468 frames from 203 unique scenes. We provide two partitions: a “Reference” partition including a paired human-edited ground truth and long-exposure shot, and a “No-reference” partition which includes handheld captures and human subjects for a more challenging evaluation (see Table 3 for details). Additional information on the test dataset such as the front and rear camera settings as well as test data examples is provided in Supp. 7.2.4
While our test set is significantly smaller than our training data, its design prioritizes perceptual relevance over sheer volume. Having a high-quality, purpose-built test set offers superior insights than a larger less controlled one, which is critical for assessing low-light image enhancement algorithms where objective metrics often fall short of human perception. Drawing inspiration from the training dataset, we categorized scenes into two broad classes: indoor and outdoor. To account for diverse usage scenarios we included both selfie and non-selfie perspectives. Moreover, we incorporated human and non-human subjects in both indoor and outdoor settings to simulate real-world conditions. Furthermore, we adhered to DXO chart capture protocols for a subset of the scenes where we controlled the capture environment to facilitate the diagnostic analysis of specific image quality attributes such as noise reduction effectiveness, accurate color rendition, and dynamic range recovery under reproducible low-light conditions. This careful dataset curation ensures that the dataset thoroughly assesses the model’s robustness and generalization across critical, perceptually distinct low-light scenarios.
@@
Figure 3: Comparison of model outputs trained on the LENVIZ and benchmark datasets. The histogram illustrates user preferences for model outputs from LENVIZ compared to others, while the spider chart highlights the advantages of various image quality aspects based on user votes. Left: LOL/SICE trained model output, Right: LENVIZ trained output
5 LENVIZ Application
To showcase the usability of our dataset for different low-light enhancement approaches, we provide an analysis on $3$ Single Exposure (LLFormer [UHDLOL2023], ExpoMamba [adhikarla2024expomamba], and ZeroDCE++ [zero_dce++]) and $3$ Multi Exposure (MEFNet [mefnet], HoLoCo [holoco], and MobileMEF [mobilemef]) low-light enhancement methods. Single exposure enhancement methods often use a low-exposure frame and attempt to improve its brightness, color, and detail recovery while reducing noise. Multi-exposure methods on the other hand incorporate the information of multiple frames captured at different exposure values, this often involves the use of a low exposure frame to recover information of over-exposed regions, and a medium or high exposure frame to get information related to color and details in under-exposed areas of the scene. All the $6$ approaches were trained using the LENVIZ dataset, as well as state-of-the-art single and multi-exposure benchmark datasets: LOL [LOL], and SICE [SICE] respectively.
5.1 Quantitative Analysis
While traditional objective metrics like PSNR are useful for quantitative reference-less assessment, they often fail to capture the perceived image quality from a human perspective, especially in low-light scenarios with complex noise and structural degradation. Taking this into account, we focused on perceptual metrics such as LPIPS and SSIM as a supplement, as these are more closely aligned with human perception. Our analysis revealed that models trained on our data consistently achieved superior scores in these metrics, demonstrating that our dataset is highly effective at training models to produce visually pleasing and structurally sound results. Furthermore, our cross-dataset evaluations confirm that these models generalize well to images from unseen cameras modules (more details on this in Supp 7.4.2). In contrast, PSNR results were more mixed, as this pixel-by-pixel metric is highly sensitive to subtle differences that might not be otherwise perceived by the human eye. A more detailed quantitative analysis, including a full table of results and cross-dataset performance, is provided in Supp 7.4. For in-depth understanding of the cases where LENVIZ trained model output we have added the analysis on failed cases in Supp 7.4.1.
5.2 Qualitative Analysis
Taking into account the limitations of quantitative metrics, we performed a comprehensive qualitative evaluation through an extensive user study (pairwise human evaluation (A/B testing)). This study provides direct insights into human perception of the enhancement quality prowess of the evaluated approaches, offering a more robust assessment in terms of visual fidelity, naturalness, and clarity under challenging illuminations. During the study, each participant was shown pairwise contrasting outputs from the LENVIZ and benchmark-trained models, participants were then asked to express their preference between the pairwise outputs and provide the primary reason for their choice. For a comprehensive evaluation, we selected 10 diverse scenes containing various objects, landscapes, illuminance levels, and both human subjects and mannequins. A total of 238 participants were randomly chosen from a pool of 89,883 Prolific candidates [^4] https://www.prolific.com/, ensuring diverse perspectives and unbiased feedback on model performance.